# Thermodynamics Reconstructed from Information Science: Theoretical Core

### v1

@@ 4.3 Mutual information and correlation free energy
At the static level, for a joint distribution $p_{XY}$ with marginals $p_{X},p_{Y}$, the mutual information
$I(X;Y):=D_{\mathrm{KL}}\!\left(p_{XY}\,\middle\|\,p_{X}p_{Y}\right)$ (25)
is the KL gap induced by the “independence projection” $p_{XY}\mapsto p_{X}p_{Y}$.
Assume the Hamiltonian is additive,
$E_{XY}(x,y)=E_{X}(x)+E_{Y}(y),$ (26)
and let $q_{X},q_{Y}$ be canonical distributions at temperature $T$:
$q_{X}(x)\propto e^{-\beta E_{X}(x)},\qquad q_{Y}(y)\propto e^{-\beta E_{Y}(y)},\qquad\beta=1/T.$ (27)
Define the nonequilibrium free-energy functionals
$F_{X}[p_{X}]:=\sum_{x}p_{X}(x)E_{X}(x)-TS(p_{X}),\qquad F_{Y}[p_{Y}]:=\sum_{y}p_{Y}(y)E_{Y}(y)-TS(p_{Y}),$ (28)
$F_{XY}[p_{XY}]:=\sum_{x,y}p_{XY}(x,y)\bigl(E_{X}(x)+E_{Y}(y)\bigr)-TS(p_{XY}),$ (29)
where $S(\cdot)$ is the Shannon entropy.
Proposition 4 (Decomposition of correlation free energy).
Under the above setting, the joint nonequilibrium free energy decomposes as
@@
Moreover, the KL divergence to the product reference $q_{X}q_{Y}$ decomposes as
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=I(X;Y)+D_{\mathrm{KL}}(p_{X}\|q_{X})+D_{\mathrm{KL}}(p_{Y}\|q_{Y}).$ (31)
Hence the “total joint nonequilibrium free energy” $TD_{\mathrm{KL}}(p_{XY}\|q_{X}q_{Y})$ decomposes exactly into subsystem nonequilibrium free energies and the correlation free energy $TI(X;Y)$.
@@
Using the entropy identity $S(p_{XY})=S(p_{X})+S(p_{Y})-I(X;Y)$,
$F_{XY}[p_{XY}]$ $=\sum_{x}p_{X}(x)E_{X}(x)+\sum_{y}p_{Y}(y)E_{Y}(y)-T\{S(p_{X})+S(p_{Y})-I(X;Y)\}$
$=F_{X}[p_{X}]+F_{Y}[p_{Y}]+TI(X;Y),$ (32)
proving (30).
Also,
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=\sum_{x,y}p_{XY}(x,y)\log\frac{p_{XY}(x,y)}{q_{X}(x)q_{Y}(y)}.$
Writing $p_{XY}(x,y)=p_{X}(x)p_{Y|X}(y|x)$ and rearranging yields
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=I(X;Y)+D_{\mathrm{KL}}(p_{X}\|q_{X})+D_{\mathrm{KL}}(p_{Y}\|q_{Y}),$
which is the static counterpart of the KL chain rule. ∎
Our framework extends beyond canonical references $q_{X},q_{Y}$ to general priors $\tau$ and Min-KL solutions with information-volume constraints, so (30) and (31) admit interpretations at a more abstract level involving measurement maps and priors.
@@ 4.4 Example: a two-state Markov-chain toy model
We consider the simplest dissipative process: a two-state Markov chain, and explicitly evaluate $\Sigma_{0,T},\Sigma_{\mathrm{obs}},\Sigma_{\mathrm{hidden}}^{(X)},\Delta S_{\mathrm{sys}},\Delta S_{\mathrm{env}},Q$.
Let the state space be $\{0,1\}$ and the transition matrix be
$P=\begin{pmatrix}1-a&a\\ b&1-b\end{pmatrix},\qquad 0<a,b<1.$ (33)
The stationary distribution is
@@
We fix the initial distribution to $\pi$.
A path on $[0,1]$ is $\omega=(x_{0},x_{1})$. The forward path measure is
$\mathbb{P}(x_{0}=i,x_{1}=j)=\pi_{i}P_{ij},$
and stationarity implies the time-reversed path measure
@@
The one-step dissipation is then
$\Sigma_{0,1}$ $=D_{\mathrm{KL}}\!\bigl(\mathbb{P}\,\big\|\,\mathbb{P}^{\mathrm{R}}\bigr)$
$=\sum_{i,j\in\{0,1\}}\pi_{i}P_{ij}\log\!\frac{\pi_{i}P_{ij}}{\pi_{j}P_{ji}}.$ (34)
This coincides with the standard expression for housekeeping dissipation in a two-state system.
We compare two extreme choices of the measurement map.
(a) Full observation.
Take $M_{\mathrm{id}}(x)=x$. Then $\Pi\mathbb{P}=\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}=\mathbb{P}^{\mathrm{R}}$, so
$\Sigma_{\mathrm{obs}}=D_{\mathrm{KL}}(\mathbb{P}\|\mathbb{P}^{\mathrm{R}})=\Sigma_{0,1},$
and $\Sigma_{\mathrm{hidden}}^{(X)}=0$.
If $\tau$ is chosen as the uniform measure on $\{0,1\}$, then $S_{M_{\mathrm{id}},\tau}$ coincides with Shannon entropy. Since both the initial and final distributions equal $\pi$,
@@
Thus the representation theorem in Sec. 5 gives
$Q=T\,\Delta S_{\mathrm{env}}=T\,\Sigma_{0,1},$
so all dissipation is interpreted as environment entropy increase (heat release).
(b) Extreme coarse-graining.
Let $M_{\mathrm{triv}}$ map both states to the same label:
$M_{\mathrm{triv}}(0)=M_{\mathrm{triv}}(1)=y_{0}.$
Then the observed path is always $(y_{0},y_{0})$, and $\Pi\mathbb{P}=\Pi\mathbb{P}^{\mathrm{R}}$, hence
$\Sigma_{\mathrm{obs}}=0,\qquad\Sigma_{\mathrm{hidden}}^{(X)}=\Sigma_{0,1}.$
Moreover, the observational entropy is identically zero (taking $\tau(\Omega)=1$), and
$\Delta S_{\mathrm{sys}}=0,\qquad\Delta S_{\mathrm{env}}=\Sigma_{0,1}.$
This example shows, in the simplest setting, that while $\Sigma_{0,1}$ is invariant, the values of observational dissipation $\Sigma_{\mathrm{obs}}$ and hidden dissipation $\Sigma_{\mathrm{hidden}}^{(X)}$ can vary drastically with the choice of observation $M$. In other words, $\Sigma_{0,T}$ is gauge-invariant, and choosing $(M,\tau)$ corresponds to a bookkeeping freedom that allocates dissipation between system and environment.
5 An information-theoretic definition of “heat”: a representation theorem from dissipation decomposition
So far we have (i) defined dissipation as the path-space KL divergence $\Sigma_{0,T}=D_{\mathrm{KL}}(\mathbb{P}\|\mathbb{P}^{\mathrm{R}})$ (Eq. (16)), and (ii) defined system entropy statically via observational entropy $S_{M,\tau}(p)$. In this section we reconstruct the conventional heat-like quantity $Q$, often introduced as “energy exchange with a bath,” as a representation theorem of dissipation decomposition that does not rely on LDB.
Define the system entropy change between times 0 and $T$ by
$\Delta S_{\mathrm{sys}}\;:=\;S_{M,\tau}(p_{T})-S_{M,\tau}(p_{0}).$ (35)
The dissipation on the same interval is, by definition,
Theorem 5 (Environment entropy change and a representation theorem for heat).
Given a measurement map $M$, a reference measure $\tau$, and path measures $\mathbb{P},\mathbb{P}^{\mathrm{R}}$, define $\Delta S_{\mathrm{sys}}$ by (35) and $\Sigma_{0,T}$ by (36). Let
$\Delta S_{\mathrm{env}}:=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}.$ (37)
Then the identity
$\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}}$ (38)
holds tautologically. Moreover, if temperature $T$ is defined as the static Min-KL dual variable via $1/T=\beta$, we may define
$Q:=T\,\Delta S_{\mathrm{env}}.$ (39)
In this way the conventional heat quantity is recovered as
$Q=T\bigl(\Sigma_{0,T}-\Delta S_{\mathrm{sys}}\bigr),$ (40)
without invoking model-dependent assumptions such as local detailed balance.
@@
By definition (37), the decomposition $\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}}$ is an identity. Since temperature $T$ is defined by $\beta=1/T$ in Sec. 3.3, multiplying (37) by $T$ and defining $Q$ by (39) yields
$Q=T\,\Sigma_{0,T}-T\,\Delta S_{\mathrm{sys}}.$
Thus $Q/T$ equals “total dissipation minus system entropy change.” ∎
@@
The theorem is an information-theoretic representation. To identify $Q$ with the actual energy exchange with a bath, one must separately specify a dynamical definition of energy and a correspondence with a bath model. The crucial point is that, even without assuming LDB or canonical bath statistics, dissipation, observational entropy, and temperature together determine a unique candidate $Q$ as a representation. In subsequent applied work, we plan to evaluate (39) for chemical reaction networks and RNAP–DNA contract-like systems, and compare with conventional definitions of heat and Landauer-type principles.
5.1 Consistency with the cross-entropy interpretation of heat
In Sec. 5, we defined total dissipation as the path relative entropy $\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$ and adopted observational entropy $S_{M,\tau}$ as system entropy. We then introduced heat $Q$ via
$\Delta S_{\mathrm{env}}:=\Sigma_{0,T}-\Delta S_{\mathrm{sys}},\qquad Q:=T\,\Delta S_{\mathrm{env}}.$ (41)
Here we clarify, conceptually, the consistency of this definition with the frequently used viewpoint that “heat (or its dimensionless form) can be expressed as a cross-entropy.”
(i) Agreement with the standard identification in isothermal environments
In standard stochastic thermodynamics (e.g. when an isothermal bath at temperature $T$ and a mechanical definition of heat $q_{\mathrm{ex}}$ are given), dissipation is written as
$\Sigma=\Delta S_{\mathrm{sys}}+\beta q_{\mathrm{ex}},\qquad\beta=\frac{1}{T}.$ (42)
Thus
$\beta q_{\mathrm{ex}}=\Sigma-\Delta S_{\mathrm{sys}}.$ (43)
Therefore (41) implies
$\frac{Q}{T}=\Delta S_{\mathrm{env}}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}.$ (44)
Hence $Q$ can be identified with the conventional (equivalent) heat $q_{\mathrm{ex}}$ in isothermal stochastic thermodynamics. The difference is conceptual: we define $Q$ from dissipation decomposition, and only later (if desired) impose model-dependent identifications such as (42).
(ii) Cross-entropy as a readable form
The cross-entropy between distributions $p$ and $q$ is
$H(p,q):=-\sum_{x}p(x)\log q(x),$ (45)
and satisfies
$H(p,q)=S(p)+D_{\mathrm{KL}}(p\|q).$ (46)
Thus, when a process naturally singles out a reference (target) distribution $p^{\star}$ and the dimensionless heat is written as
$\frac{Q}{T}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}},$ (47)
if the right-hand side can be evaluated/rearranged to contain a KL term $D_{\mathrm{KL}}(\cdot\|p^{\star})$, then (46) enables a rewriting of $Q/T$ in terms of a cross-entropy $H(\cdot,p^{\star})$. In this sense, (47) acts as a master bookkeeping identity that contains cross-entropy expressions as special readable forms.
Typical situations include:
  1. (a)
    Relaxation/erasure processes toward a target $p^{\star}$ in which dissipation (or its dominant part) reduces to a form like $D_{\mathrm{KL}}(p_{0}\|p^{\star})$. Then (46) produces $S(p_{0})+D_{\mathrm{KL}}(p_{0}\|p^{\star})=H(p_{0},p^{\star})$, allowing $Q/T$ to be interpreted as a cross-entropy.
  2. (b)
    Processes where the final state becomes (in a coarse-grained sense) sharply localized so that $S(p_{T})\simeq 0$ (e.g. logical erasure). Then $\Delta S_{\mathrm{sys}}$ is effectively fixed, and if $\Sigma_{0,T}$ contains a term $D_{\mathrm{KL}}(p_{0}\|p^{\star})$, $Q/T$ tends to simplify to $H(p_{0},p^{\star})$ (up to state-function terms).
(iii) Connection to Landauer-type arguments in Tsuruyama (2025)
Tsuruyama [17] starts from a decomposition of the form $\sigma=\Delta S+\beta q_{\mathrm{ex}}$ in a Landauer-type erasure setting and rearranges $\beta q_{\mathrm{ex}}$ to obtain, for a binary erasure problem,
$\beta q_{\mathrm{ex}}=-p\log p^{\ast}-(1-p)\log(1-p^{\ast}),$ (48)
which is a cross-entropy form. This is naturally interpreted as a concrete realization of
$\beta q_{\mathrm{ex}}=\Sigma-\Delta S_{\mathrm{sys}}$ (49)
via (46). Therefore, our representation (41) contains Tsuruyama’s “heat = cross-entropy” redefinition as a special case, without assuming a specific model class or erasure protocol at the level of definitions.
(iv) Implications of using observational entropy $S_{M,\tau}$
We define $\Delta S_{\mathrm{sys}}$ using observational entropy $S_{M,\tau}$ rather than Shannon entropy. This does not break consistency with cross-entropy interpretations. Instead, because $S_{M,\tau}$ explicitly includes coarse-graining cell volumes $\tau(C_{y})$, it naturally captures situations where “the logical state is fixed but microstate volume (information volume) remains.” When one rewrites $Q/T$ in cross-entropy form, the essential choice is the reference (target) distribution $p^{\star}$; in our language, this corresponds to the gauge choice of the measurement map $M$ and the reference measure $\tau$ (the descriptive convention).

### v2

@@ 4.3 Mutual information and correlation free energy
For a static joint distribution $p_{XY}$ with marginals $p_{X},p_{Y}$, mutual information is
$I(X;Y)=D_{\mathrm{KL}}(p_{XY}\|p_{X}p_{Y}).$ (23)
Assume the Hamiltonian is additive:
$E_{XY}(x,y)=E_{X}(x)+E_{Y}(y).$ (24)
Let $q_{X},q_{Y}$ be canonical references at temperature $T$:
$q_{X}(x)\propto e^{-\beta E_{X}(x)},\qquad q_{Y}(y)\propto e^{-\beta E_{Y}(y)}.$ (25)
Define nonequilibrium free energies:
$F_{X}[p_{X}]=\sum_{x}p_{X}(x)E_{X}(x)-TS(p_{X}),\qquad F_{Y}[p_{Y}]=\sum_{y}p_{Y}(y)E_{Y}(y)-TS(p_{Y}),$
$F_{XY}[p_{XY}]=\sum_{x,y}p_{XY}(x,y)(E_{X}(x)+E_{Y}(y))-TS(p_{XY}).$
Proposition 3 (Free-energy decomposition with correlation term).
Under the above assumptions,
@@
Moreover,
$D_{\mathrm{KL}}(p_{XY}\|q_{X}q_{Y})=I(X;Y)+D_{\mathrm{KL}}(p_{X}\|q_{X})+D_{\mathrm{KL}}(p_{Y}\|q_{Y}).$ (27)
@@
Use the entropy identity $S(p_{XY})=S(p_{X})+S(p_{Y})-I(X;Y)$ to obtain (26). The KL decomposition (27) follows by writing $p_{XY}=p_{X}p_{Y|X}$ and rearranging terms. ∎
@@ 4.4 Example: a two-state Markov-chain toy model
We illustrate the general theory using the simplest irreversible stochastic dynamics: a two-state, time-homogeneous Markov chain. Despite its minimal structure, this model already displays all key features of path-space dissipation, its dependence on observational resolution, and the decomposition into observed and hidden components.
Let the state space be $\{0,1\}$, and let the transition matrix be
$P=\begin{pmatrix}1-a&a\\ b&1-b\end{pmatrix},\qquad 0<a,b<1,$ (28)
so that both states communicate. The stationary distribution is
@@
A one-step trajectory is $(x_{0},x_{1})$ drawn from $\mathbb{P}(x_{0}=i,x_{1}=j)=\pi_{i}P_{ij}$. Stationarity implies that the corresponding time-reversed one-step trajectory distribution is
@@
Thus the dissipation (entropy production) over one step is the path-space KL divergence
$\Sigma_{0,1}=\sum_{i,j\in\{0,1\}}\pi_{i}P_{ij}\,\log\!\frac{\pi_{i}P_{ij}}{\pi_{j}P_{ji}},$ (29)
which is strictly positive unless detailed balance holds ($\pi_{i}P_{ij}=\pi_{j}P_{ji}$). This quantity provides the benchmark against which the effects of coarse-graining will be measured.
4.5 Full observation
4.5.1 (a) Full observation $M_{\mathrm{id}}$.
Under full observation, each microstate transition is visible, so the observational channel is the identity. Therefore $\Pi\mathbb{P}=\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}=\mathbb{P}^{\mathrm{R}}$, giving
$\Sigma_{\mathrm{obs}}=\Sigma_{0,1},\qquad\Sigma_{\mathrm{hidden}}^{(X)}=0.$
If the reference measure is uniform on $\{0,1\}$, the observational entropy $S_{M_{\mathrm{id}},\tau}$ coincides with Shannon entropy. Because the system starts and ends in the same stationary distribution, the system entropy change is zero:
$\Delta S_{\mathrm{sys}}=0.$
Therefore all dissipation is allocated to the environment:
$\Delta S_{\mathrm{env}}=\Sigma_{0,1},\qquad Q=T\,\Sigma_{0,1}.$
This case shows that in a fully resolved description, the dissipation is entirely revealed at the observational level; no hidden dissipation remains.
(b) Extreme coarse-graining $M_{\mathrm{triv}}$.
Consider the opposite observational limit, in which the measurement map collapses both states into a single outcome. Let $M_{\mathrm{triv}}(0)=M_{\mathrm{triv}}(1)=y_{0}$. Then every path is mapped to the same constant trajectory $(y_{0},y_{0})$ both forward and backward, so
$\Pi\mathbb{P}=\Pi\mathbb{P}^{\mathrm{R}},\qquad\Sigma_{\mathrm{obs}}=0.$
Thus the entire dissipation becomes invisible at the observational scale:
$\Sigma_{\mathrm{hidden}}^{(X)}=\Sigma_{0,1}.$
With a single observational cell, the observational entropy is constant (and may be taken as zero), so again
@@ 4.5 Full observation
Unlike the fully observed case, here the system appears perfectly reversible even though the underlying dynamics is irreversible. All dissipation is attributed to the “environment” due solely to loss of resolution in $M$. All dissipation is attributed to the “environment” due solely to loss of resolution in $M$. This explicitly demonstrates that the system–environment split in entropy accounting is gauge-dependent, whereas the total dissipation $\Sigma_{0,1}$ is invariant. This toy model highlights three essential principles of the general theory:
  1. 1.
    Even the simplest Markov chain exhibits nonzero dissipation whenever detailed balance is broken.
  2. 2.
    Observational resolution controls how much dissipation is observable; coarse-graining always decreases the observed part, but the total dissipation remains invariant.
  3. 3.
    The decomposition of dissipation into system and environment terms depends on the chosen observational gauge, whereas the KL divergence on path space provides the unique, gauge-invariant notion of irreversibility.
These features extend directly to high-dimensional or continuous-state dynamics, where coarse-graining may hide massive amounts of dissipation, making the KL projection-gap interpretation crucial for analyzing nonequilibrium systems.
5 Heat as a representation theorem
We now combine the static intensive variable $T$ with the dynamic dissipation $\Sigma_{0,T}$. Because $\Sigma_{0,T}$ is model-independent and uniquely defined by path-time asymmetry, and because the system entropy $S_{M,\tau}$ is fixed by the choice of $(M,\tau)$, it follows that the only consistent definition of heat must be constructed from these two quantities.
5.1 Definition of heat
Let a process start from $p_{0}$ and end at $p_{T}$. We have the decomposition
$\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}},\qquad\Delta S_{\mathrm{sys}}:=S_{M,\tau}(p_{T})-S_{M,\tau}(p_{0}).$ (30)
Definition 1 (Heat).
Define
$Q:=T\,\Delta S_{\mathrm{env}}=T\bigl(\Sigma_{0,T}-\Delta S_{\mathrm{sys}}\bigr).$ (31)
Thus heat is determined uniquely by the pair $(T,\Sigma_{0,T})$ once the gauge $(M,\tau)$ has been fixed.
Remark 2 (Connection to standard stochastic-thermodynamics heat under LDB).
In conventional stochastic thermodynamics, one often defines a trajectory-level heat $Q_{\mathrm{LDB}}[\gamma]$ via energy exchange with an explicit bath, assuming local detailed balance (LDB). In that setting the path-probability ratio admits the form
$\log\frac{\,\mathrm{d}\mathbb{P}[\gamma]}{\,\mathrm{d}\mathbb{P}^{\mathrm{R}}[\gamma^{\mathrm{R}}]}=\Delta s_{\mathrm{sys}}[\gamma]+\beta\,Q_{\mathrm{LDB}}[\gamma],$
so that upon averaging one obtains $\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\beta\,\langle Q_{\mathrm{LDB}}\rangle$. Therefore, when $(M,\tau)$ is chosen so that $S_{M,\tau}$ reduces to the usual Gibbs–Shannon entropy, our definition $Q:=T(\Sigma_{0,T}-\Delta S_{\mathrm{sys}})$ coincides with the standard mean heat $\langle Q_{\mathrm{LDB}}\rangle$. The present framework keeps this consistency while not requiring LDB or an explicit bath model: Eq. (31) defines the unique heat scalar compatible with (i) the Min-KL temperature, (ii) path-KL dissipation as irreversibility, and (iii) the chosen entropic bookkeeping $S_{M,\tau}$.
Proposition 4 (Representation theorem for heat).
Heat is uniquely determined by the following axioms:
  1. (i)
    (Static consistency) For quasistatic processes connecting Min-KL equilibrium states, $Q=T\,\Delta S_{\mathrm{env}}$ reproduces the standard thermodynamic relation.
  2. (ii)
    (Dynamic consistency) For arbitrary processes, irreversibility is quantified by $\Sigma_{0,T}$.
  3. (iii)
    (Gauge covariance) System entropy depends on $(M,\tau)$ as $S_{M,\tau}$; dissipation $\Sigma_{0,T}$ is invariant; heat transforms as a scalar $Q=T(\Sigma_{0,T}-\Delta S_{\mathrm{sys}})$.
Then $Q$ must be given by (31).
@@ 5.1 Definition of heat
Given the decomposition (30), the only allowed scalar combination of $(T,\Sigma_{0,T},\Delta S_{\mathrm{sys}})$ that satisfies (i)–(iii) is $Q=T(\Sigma_{0,T}-\Delta S_{\mathrm{sys}})$. Any alternative expression would violate either static consistency (quasistatic limit), dynamic consistency (monotonicity with dissipation), or gauge covariance. ∎
5.2 Cross-entropy consistency and entropy flows
We now justify (31) by a cross-entropy identity. Let $p_{0},p_{T}$ be the initial and final distributions, and $p^{\ast}_{0},p^{\ast}_{T}$ the corresponding Min-KL equilibrium states with respect to $(M,\tau)$.
The cross-entropy $\mathrm{CE}(p\|q):=-\int p\log q$ satisfies
$\mathrm{CE}(p_{T}\|p_{T}^{\ast})-\mathrm{CE}(p_{0}\|p_{0}^{\ast})=D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)_{T}\|p_{T}^{\ast})-D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)_{0}\|p_{0}^{\ast})+S_{M,\tau}(p_{T})-S_{M,\tau}(p_{0}).$
Using that $p^{\ast}_{t}$ satisfies the thermodynamic relation $U_{t}=U[p^{\ast}_{t}]$, one obtains
$\Delta S_{\mathrm{sys}}+\frac{Q}{T}=\Sigma_{0,T}+\bigl[D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)_{T}\|p_{T}^{\ast})-D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)_{0}\|p_{0}^{\ast})\bigr].$ (32)
For processes that start or end near equilibrium, the bracketed term is $o(1)$, so the identity reduces to (31). Thus the proposed heat expression is cross-entropy consistent.
@@ 5.2 Cross-entropy consistency and entropy flows
Equation (32) also clarifies the role of nonequilibrium free energy:
$F[p]=U[p]-TS_{M,\tau}(p)=F[p^{\ast}]+T\,D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)\|p^{\ast}).$
Changes in nonequilibrium free energy appear as the KL terms in (32).